OpenStudy (anonymous):

√x - 7 + 5 = 11 show work

OpenStudy (anonymous):

Is that \[\sqrt{x-7}+5=11\] or \[\sqrt{x}-7+5=11\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):